# Expert-Data Alignment Governs Generation Quality in Decentralized Diffusion Models

### v1

@@ 1 Introduction
Decentralized Diffusion Models (DDMs) (McAllister et al., 2025) combine independently trained diffusion experts (Ho et al., 2020) via an inference-time router. Because experts are trained on disjoint data clusters and can strongly disagree in their predictions, understanding what governs generation quality becomes crucial—yet this question has not been systematically studied.
A natural hypothesis is that numerical stability determines quality (Yang et al., 2023), that is, routing strategies that minimize trajectory sensitivity should produce superior samples. We demonstrate that this hypothesis is incorrect. In this work, we observe that full ensemble routing, in which all expert predictions are combined, achieves the lowest trajectory sensitivity and best numerical convergence, yet produces the worst generation quality. These results rule out trajectory sensitivity as the primary determinant of generation quality.
@@
Although numerical stability does not govern quality, understanding when DDM sampling converges remains valuable. Classical stability analysis suggests DDMs should fail: Lipschitz constants of deep networks grow exponentially with depth (Fazlyab et al., 2019; Virmaux and Scaman, 2018; Yang et al., 2023), and Grönwall’s inequality implies small perturbations amplify over integration (Hairer et al., 1993). We also examine trajectory-local sensitivity, denoted $\widehat{L}_{\text{eff}}^{(h)}$, which formalizes the idea that Jacobian spectral norms remain bounded along realized sampling paths. Empirically, we observe that trajectories exhibit moderate sensitivity compared to worst-case global bounds. While $\widehat{L}_{\text{eff}}^{(h)}$ does not predict quality across routing strategies, it may serve as a within-strategy diagnostic for identifying numerically sensitive samples.
@@
  • Stability–quality dissociation. We demonstrate that trajectory sensitivity does not govern generation quality: full ensemble routing achieves the lowest $\widehat{L}_{\text{eff}}^{(h)}$ and step-refinement disagreement, yet produces the worst Fréchet Inception Distance (FID) (Heusel et al., 2017). This rules out numerical stability as the primary quality determinant. We additionally explore trajectory-local sensitivity as a within-strategy diagnostic, finding weak predictive power.
@@ 2 Related Work
Diffusion sampling involves Lipschitz-constrained ODEs where discretization error affects accuracy. Sampling can be expressed as a probability-flow Ordinary Differential Equation (ODE) (Song et al., 2020), where Lipschitz constants and discretization error determine solver accuracy. Recent work by Tan et al. (2025) addresses temporal stiffness via a specialized solver called STORK based on stabilized Runge-Kutta methods for stiff diffusion ODEs. STORK’s stabilized solvers could, in principle, be combined with our approach (using stable solvers for temporal stiffness while using sparse routing to control spatial sensitivity). We use the released Euler and Heun solvers to isolate the effect of routing strategies, but STORK-style solvers may provide additional gains in decentralized settings.
@@
Local Lipschitz bounds have been extensively studied in the neural network literature. Jordan and Dimakis (2020) developed methods for exactly computing local Lipschitz constants, enabling input-specific sensitivity analysis rather than global worst-case bounds. Our work applies this trajectory-local perspective to diffusion sampling. The key novelty is not the local Lipschitz concept itself, but its application to understanding decentralized diffusion dynamics and the discovery that routing implicitly stabilizes the sampling dynamics.
@@
DDMs show that decentralized experts, when routed, can match a monolithic diffusion objective (McAllister et al., 2025). Our work investigates why such combination succeeds despite expert disagreement, providing a stability-based explanation complementary to the original capacity arguments.
@@
DDMs are ensembles of independently trained models, not traditional Mixture-of-Experts (MoE). In standard MoE architectures experts are FFN layers within a shared backbone, trained jointly with load balancing losses, and routed at the token level (Shazeer et al., 2017; Fedus et al., 2022). In DDM, each “expert” is a complete diffusion model trained in isolation on a disjoint data partition (no shared parameters, no gradient communication, no joint training). Routing occurs at the input level (entire noisy images) rather than token level, and experts are combined only at inference time. Concurrent work applies traditional MoE architectures within diffusion models (Fei et al., 2024; Sun et al., 2024; Shi et al., 2025; Yuan et al., 2025; Cheng et al., 2025; Liu et al., 2025; Zheng et al., 2025); DDM instead combines complete, independently trained models. This distinction matters for stability analysis: DDM experts can produce arbitrarily different outputs for the same input (having never coordinated during training), whereas MoE experts share a representational backbone that constrains their disagreement. Our analysis specifically addresses the stability challenges arising from DDM’s decentralized training.
@@
The importance of matching inputs to appropriately trained experts is well-established in MoE systems. Sparsely-gated MoE architectures rely on learned routing to direct inputs to relevant experts (Shazeer et al., 2017), with subsequent work analyzing expert utilization and load balancing (Dai et al., 2022; Zhou et al., 2022). In federated learning, data heterogeneity across clients creates analogous challenges: models trained on non-IID partitions may produce poor predictions on out-of-distribution inputs (Li et al., 2020). Our work provides direct experimental evidence that this principle governs sample quality in DDMs: sparse routing succeeds precisely because it maintains alignment between inputs and expert training distributions.
@@ 3.2 Diffusion sampling
Diffusion sampling generates data by integrating an ordinary differential equation (ODE) $\frac{dx_{t}}{dt}=v_{t}(x_{t})$ from noise to data (Lipman et al., 2022). We call $v=\{v_{t}\}_{t=0}^{1}$ a flow and the solution $\{x_{t}\}$ a trajectory. The ODE has a unique solution when $v$ is Lipschitz continuous (Coddington et al., 1956).
@@ 3.3 Decentralized Diffusion
Decentralized Diffusion Models (DDMs) (McAllister et al., 2025) train $K$ expert diffusion models in complete isolation on disjoint data partitions. Unlike traditional MoE (which routes tokens to different FFN layers within a shared backbone), DDM routes entire inputs to separate full models at each denoising step.
@@
The router outputs probabilities $p_{t}(k|x_{t})$ for each expert $k$. Full ensemble mode uses $w_{t}^{(k)}(x_{t})=p_{t}(k|x_{t})$ for all $k$. Top-$k$ routing selects the $k$ experts with highest probability and renormalizes their weights (Shazeer et al., 2017). Top-1 selects a single expert: $v_{\text{Top-1}}(x_{t})=v_{k^{*}}(x_{t})$ with $k^{*}=\arg\max_{k}p_{t}(k|x_{t})$. This yields a piecewise-smooth vector field with non-differentiability on measure-zero switching surfaces.
@@ 4 The Stability–Quality Dissociation
Sections 5 and 6 will demonstrate the dissociation: full ensemble routing achieves the lowest trajectory sensitivity and lowest step-refinement disagreement, yet produces the worst generation quality—even worse generation quality is observed in (Jiang et al., 2025). Top-2 achieves the best generation quality despite higher trajectory sensitivity. This rules out numerical stability as the quality determinant.
@@ 5.2 Cluster Distance Analysis
We use the pretrained DDM Paris model (Jiang et al., 2025) via released pretrained checkpoints. The model consists of $K=8$ experts trained on a subset of LAION-Aesthetics (Schuhmann et al., 2022). The dataset was partitioned into 8 semantic clusters via two-stage hierarchical k-means on DINOv2-ViT-L/14 embeddings. The model uses a DiT-B/2 router ($\sim$129M parameters) and 8 DiT-XL/2 experts (a modified version of the DiT-XL/2 experts with $\sim$606M parameters each, $\sim$5B total). The router was trained post-hoc on the full dataset, effectively learning to route inputs to the expert trained on the most similar data. We will use Paris DDM to test the Expert-Data Alignment Hypothesis.
@@ 5.4 Expert Disagreement and Sample Quality
For $n=500$ samples generated with full ensemble, we measure (1) mean pairwise expert disagreement $D(x_{t})=\frac{1}{\binom{K}{2}}\sum_{i<j}\|v_{i}(x_{t})-v_{j}(x_{t})\|_{2}$, (2) trajectory-integrated disagreement: $D_{\text{int}}=\int_{0}^{1}D(x_{t})\,dt$, and (3) perceptual quality as LPIPS distance (Zhang et al., 2018) to the corresponding Top-2 output (matched initial noise). See Figure 1 for the results.
@@ 6.2.2 Experiment 2: Measuring Trajectory-Local Sensitivity
Table 4: Routing strategy comparison. FID from (Jiang et al., 2025); $\widehat{L}_{\text{eff}}^{(h)}$ and $\Delta_{\text{refine}}$ measured on n=1000n{=}1000 samples. Full ensemble achieves the lowest trajectory sensitivity and step-refinement disagreement yet produces the worst FID.
@@ 6.2.3 Experiment 3: Step-Size Refinement and Convergence
where $D(\cdot)$ denotes VAE decoding, LPIPS measures perceptual distance in decoded image space (Zhang et al., 2018), and $x_{1}$ is the initial noise. The metric $\Delta_{\text{refine}}$ is an a posteriori refinement-based error indicator that does not involve Jacobian computation, providing a measure of numerical integration error that avoids methodological coupling with $L_{\text{eff}}$.
@@ Appendix B Probabilistic Convergence under Empirical Stability
Step 1 (Deterministic bound on $A_{L}$). Standard Grönwall-based global error analysis for one-step methods (see, e.g., Hairer et al. (1993)) shows that if the velocity field has Lipschitz constant $L$ along the trajectory, then the global error satisfies
@@ Appendix C Expert-Data Alignment Experiment Details
We use DINOv2-ViT-L/14 (Oquab et al., 2023) to extract embeddings for both training data cluster centroids and intermediate sampling states. For intermediate states $x_{t}$ during sampling, we first decode through the VAE to obtain pixel-space images, then extract DINOv2 embeddings. The 8 cluster centroids were computed during DDM training using hierarchical k-means on DINOv2 embeddings of the training set.
@@ D.1 Experimental Setup
We use the pretrained DDM Paris model (Jiang et al., 2025) via released pretrained checkpoints. The model consists of $K=8$ experts trained on a subset of LAION-Aesthetics (Schuhmann et al., 2022). The dataset was partitioned into 8 semantic clusters via two-stage hierarchical k-means on DINOv2-ViT-L/14 embeddings. The model uses a DiT-B/2 router ($\sim$129M parameters) and 8 DiT-XL/2 experts (Peebles and Xie, 2023) (a modified version of the DiT-XL/2 experts with $\sim$606M parameters each, $\sim$5B total).
We compare Top-1, Top-2, and full-ensemble routing strategies on the DDM architecture (McAllister et al., 2025). To avoid circularity, that is defining failure via $L_{\text{eff}}$ and then claiming $L_{\text{eff}}$ predicts failure, we use Eq.(2) as a methodologically independent error signal.
@@ D.2 Experiment 1: Local Truncation Error
Following Section 6.2.1, we estimate one-step local truncation error by comparing a single Heun step of size $h$ (Karras et al., 2022) against a higher-precision reference obtained by subdividing the same interval into 10 Heun sub-steps of size $h/10$. For each routing strategy, we sample 1,000 trajectories and evaluate at randomly selected trajectory points $(x_{t},t)$ the maximum local truncation error defined as $\epsilon^{\text{local}}=\max_{n}\|\epsilon_{n}^{\text{local}}(t_{n})\|=Ch^{p+1}$ for some constant $C$. Table 5 reports these measurements.
@@ F.5 Generalization Tests Without Retraining
We evaluate the fixed trained experts and router on COCO captions as prompts. We sample n=100n{=}100 prompts uniformly at random from the caption set, generate one sample per prompt with fixed seeds, and compute $\widehat{L}_{\text{eff}}^{(h)}$ and $\Delta_{\text{refine}}$ as in the main text. We additionally report the AUC of $\widehat{L}_{\text{eff}}^{(h)}$ for predicting high $\Delta_{\text{refine}}$ events using the same percentile thresholding protocol. We repeat the same evaluation under a harder numerical regime without retraining: (i) fewer solver steps (Heun-25 instead of Heun-50), and (ii) higher CFG (Ho and Salimans, 2022) (7.5 instead of 4.0). The goal is not quality, but whether sensitivity rankings and predictiveness persist. Table 14 compares baseline and stressed regimes.

### v2

@@ 1 Introduction
Decentralized Diffusion Models (DDMs) [mcallister2025ddm] combine independently trained diffusion experts [ho2020denoising] via an inference-time router. Because experts are trained on disjoint data clusters and can strongly disagree in their predictions, understanding what governs generation quality becomes crucial—yet this question has not been systematically studied.
A natural hypothesis is that numerical stability determines quality [yang2024lipschitz], that is, routing strategies that minimize trajectory sensitivity should produce superior samples. We demonstrate that this hypothesis is incorrect. In this work, we observe that full ensemble routing, in which all expert predictions are combined, achieves the lowest trajectory sensitivity and best numerical convergence, yet produces the worst generation quality. These results rule out trajectory sensitivity as the primary determinant of generation quality.
@@
Although numerical stability does not govern quality, understanding when DDM sampling converges remains valuable. Classical stability analysis suggests DDMs should fail: Lipschitz constants of deep networks grow exponentially with depth [fazlyab2019efficient, virmaux2018lipschitz, yang2024lipschitz], and Grönwall’s inequality implies small perturbations amplify over integration [hairer1993solving]. We also examine trajectory-local sensitivity, denoted $\widehat{L}_{\text{eff}}^{(h)}$, which formalizes the idea that Jacobian spectral norms remain bounded along realized sampling paths. Empirically, we observe that trajectories exhibit moderate sensitivity compared to worst-case global bounds. While $\widehat{L}_{\text{eff}}^{(h)}$ does not predict quality across routing strategies, it may serve as a within-strategy diagnostic for identifying numerically sensitive samples.
@@
  • Stability–quality dissociation. We demonstrate that trajectory sensitivity does not govern generation quality: full ensemble routing achieves the lowest $\widehat{L}_{\text{eff}}^{(h)}$ and step-refinement disagreement, yet produces the worst Fréchet Inception Distance (FID) [heusel2017gans]. This rules out numerical stability as the primary quality determinant. We additionally explore trajectory-local sensitivity as a within-strategy diagnostic, finding weak predictive power.
@@ 2 Related Work
Diffusion sampling involves Lipschitz-constrained ODEs where discretization error affects accuracy. Sampling can be expressed as a probability-flow Ordinary Differential Equation (ODE) [song2021scorebased], where Lipschitz constants and discretization error determine solver accuracy. Recent work by tan2025stork addresses temporal stiffness via a specialized solver called STORK based on stabilized Runge-Kutta methods for stiff diffusion ODEs. STORK’s stabilized solvers could, in principle, be combined with our approach (using stable solvers for temporal stiffness while using sparse routing to control spatial sensitivity). We use the released Euler and Heun solvers to isolate the effect of routing strategies, but STORK-style solvers may provide additional gains in decentralized settings.
@@
Local Lipschitz bounds have been extensively studied in the neural network literature. jordan2020exactly developed methods for exactly computing local Lipschitz constants, enabling input-specific sensitivity analysis rather than global worst-case bounds. Our work applies this trajectory-local perspective to diffusion sampling. The key novelty is not the local Lipschitz concept itself, but its application to understanding decentralized diffusion dynamics and the discovery that routing implicitly stabilizes the sampling dynamics.
@@
DDMs show that decentralized experts, when routed, can match a monolithic diffusion objective [mcallister2025ddm]. Our work investigates why such combination succeeds despite expert disagreement, providing a stability-based explanation complementary to the original capacity arguments.
@@
DDMs are ensembles of independently trained models, not traditional Mixture-of-Experts (MoE). In standard MoE architectures experts are FFN layers within a shared backbone, trained jointly with load balancing losses, and routed at the token level [shazeer2017outrageously, fedus2021switch]. In DDM, each “expert” is a complete diffusion model trained in isolation on a disjoint data partition (no shared parameters, no gradient communication, no joint training). Routing occurs at the input level (entire noisy images) rather than token level, and experts are combined only at inference time. Concurrent work applies traditional MoE architectures within diffusion models [fei2024ditmoe, sun2025ecdit, yang2025diffmoe, wang2025expertrace, cheng2025diffmoe, liu2025dsmoe, zheng2025dense2moe]; DDM instead combines complete, independently trained models. This distinction matters for stability analysis: DDM experts can produce arbitrarily different outputs for the same input (having never coordinated during training), whereas MoE experts share a representational backbone that constrains their disagreement. Our analysis specifically addresses the stability challenges arising from DDM’s decentralized training.
@@
The importance of matching inputs to appropriately trained experts is well-established in MoE systems. Sparsely-gated MoE architectures rely on learned routing to direct inputs to relevant experts [shazeer2017outrageously], with subsequent work analyzing expert utilization and load balancing [dai2022stablemoe, zhou2022expertchoice]. In federated learning, data heterogeneity across clients creates analogous challenges: models trained on non-IID partitions may produce poor predictions on out-of-distribution inputs [li2020federated]. Our work provides direct experimental evidence that this principle governs sample quality in DDMs: sparse routing succeeds precisely because it maintains alignment between inputs and expert training distributions.
@@ 3.2 Diffusion sampling
Diffusion sampling generates data by integrating an ordinary differential equation (ODE) $\frac{dx_{t}}{dt}=v_{t}(x_{t})$ from noise to data [lipman2023flow]. We call $v=\{v_{t}\}_{t=0}^{1}$ a flow and the solution $\{x_{t}\}$ a trajectory. The ODE has a unique solution when $v$ is Lipschitz continuous [coddington1955theory].
@@ 3.3 Decentralized Diffusion
Decentralized Diffusion Models (DDMs) [mcallister2025ddm] train $K$ expert diffusion models in complete isolation on disjoint data partitions. Unlike traditional MoE (which routes tokens to different FFN layers within a shared backbone), DDM routes entire inputs to separate full models at each denoising step.
@@
The router outputs probabilities $p_{t}(k|x_{t})$ for each expert $k$. Full ensemble mode uses $w_{t}^{(k)}(x_{t})=p_{t}(k|x_{t})$ for all $k$. Top-$k$ routing selects the $k$ experts with highest probability and renormalizes their weights [shazeer2017outrageously]. Top-1 selects a single expert: $v_{\text{Top-1}}(x_{t})=v_{k^{*}}(x_{t})$ with $k^{*}=\arg\max_{k}p_{t}(k|x_{t})$. This yields a piecewise-smooth vector field with non-differentiability on measure-zero switching surfaces.
@@ 4 The Stability–Quality Dissociation
Sections 5 and 6 will demonstrate the dissociation: full ensemble routing achieves the lowest trajectory sensitivity and lowest step-refinement disagreement, yet produces the worst generation quality—even worse generation quality is observed in [jiang2025paris]. Top-2 achieves the best generation quality despite higher trajectory sensitivity. This rules out numerical stability as the quality determinant.
@@ 5.2 Cluster Distance Analysis
We use the pretrained DDM Paris model [jiang2025paris] via released pretrained checkpoints. The model consists of $K=8$ experts trained on a subset of LAION-Aesthetics [schuhmann2022laion]. The dataset was partitioned into 8 semantic clusters via two-stage hierarchical k-means on DINOv2-ViT-L/14 embeddings. The model uses a DiT-B/2 router ($\sim$129M parameters) and 8 DiT-XL/2 experts (a modified version of the DiT-XL/2 experts with $\sim$606M parameters each, $\sim$5B total). The router was trained post-hoc on the full dataset, effectively learning to route inputs to the expert trained on the most similar data. We will use Paris DDM to test the Expert-Data Alignment Hypothesis.
@@ 5.4 Expert Disagreement and Sample Quality
For $n=500$ samples generated with full ensemble, we measure (1) mean pairwise expert disagreement $D(x_{t})=\frac{1}{\binom{K}{2}}\sum_{i<j}\|v_{i}(x_{t})-v_{j}(x_{t})\|_{2}$, (2) trajectory-integrated disagreement: $D_{\text{int}}=\int_{0}^{1}D(x_{t})\,dt$, and (3) perceptual quality as LPIPS distance [zhang2018unreasonable] to the corresponding Top-2 output (matched initial noise). See Figure 1 for the results.
@@ 6.2.2 Experiment 2: Measuring Trajectory-Local Sensitivity
Table 4: Routing strategy comparison. FID from [jiang2025paris]; $\widehat{L}_{\text{eff}}^{(h)}$ and $\Delta_{\text{refine}}$ measured on n=1000n{=}1000 samples. Full ensemble achieves the lowest trajectory sensitivity and step-refinement disagreement yet produces the worst FID.
@@ 6.2.3 Experiment 3: Step-Size Refinement and Convergence
where $D(\cdot)$ denotes VAE decoding, LPIPS measures perceptual distance in decoded image space [zhang2018unreasonable], and $x_{1}$ is the initial noise. The metric $\Delta_{\text{refine}}$ is an a posteriori refinement-based error indicator that does not involve Jacobian computation, providing a measure of numerical integration error that avoids methodological coupling with $L_{\text{eff}}$.
@@ Appendix B Probabilistic Convergence under Empirical Stability
Step 1 (Deterministic bound on $A_{L}$). Standard Grönwall-based global error analysis for one-step methods (see, e.g., hairer1993solving) shows that if the velocity field has Lipschitz constant $L$ along the trajectory, then the global error satisfies
@@ Appendix C Expert-Data Alignment Experiment Details
We use DINOv2-ViT-L/14 [oquab2023dinov2] to extract embeddings for both training data cluster centroids and intermediate sampling states. For intermediate states $x_{t}$ during sampling, we first decode through the VAE to obtain pixel-space images, then extract DINOv2 embeddings. The 8 cluster centroids were computed during DDM training using hierarchical k-means on DINOv2 embeddings of the training set.
@@ D.1 Experimental Setup
We use the pretrained DDM Paris model [jiang2025paris] via released pretrained checkpoints. The model consists of $K=8$ experts trained on a subset of LAION-Aesthetics [schuhmann2022laion]. The dataset was partitioned into 8 semantic clusters via two-stage hierarchical k-means on DINOv2-ViT-L/14 embeddings. The model uses a DiT-B/2 router ($\sim$129M parameters) and 8 DiT-XL/2 experts [peebles2023dit] (a modified version of the DiT-XL/2 experts with $\sim$606M parameters each, $\sim$5B total).
We compare Top-1, Top-2, and full-ensemble routing strategies on the DDM architecture [mcallister2025ddm]. To avoid circularity, that is defining failure via $L_{\text{eff}}$ and then claiming $L_{\text{eff}}$ predicts failure, we use Eq.(2) as a methodologically independent error signal.
@@ D.2 Experiment 1: Local Truncation Error
Following Section 6.2.1, we estimate one-step local truncation error by comparing a single Heun step of size $h$ [karras2022elucidating] against a higher-precision reference obtained by subdividing the same interval into 10 Heun sub-steps of size $h/10$. For each routing strategy, we sample 1,000 trajectories and evaluate at randomly selected trajectory points $(x_{t},t)$ the maximum local truncation error defined as $\epsilon^{\text{local}}=\max_{n}\|\epsilon_{n}^{\text{local}}(t_{n})\|=Ch^{p+1}$ for some constant $C$. Table 5 reports these measurements.
@@ F.5 Generalization Tests Without Retraining
We evaluate the fixed trained experts and router on COCO captions as prompts. We sample n=100n{=}100 prompts uniformly at random from the caption set, generate one sample per prompt with fixed seeds, and compute $\widehat{L}_{\text{eff}}^{(h)}$ and $\Delta_{\text{refine}}$ as in the main text. We additionally report the AUC of $\widehat{L}_{\text{eff}}^{(h)}$ for predicting high $\Delta_{\text{refine}}$ events using the same percentile thresholding protocol. We repeat the same evaluation under a harder numerical regime without retraining: (i) fewer solver steps (Heun-25 instead of Heun-50), and (ii) higher CFG [ho2022classifier] (7.5 instead of 4.0). The goal is not quality, but whether sensitivity rankings and predictiveness persist. Table 14 compares baseline and stressed regimes.